give an example of three f,g, and h (none of which is a constant function) such that fog=foh, but g is not equal to h.
Answer by richard1234(5244)

Let

f(x) = x^2
g(x) = x
h(x) = -x

Then f(g(x)) = x^2 and f(h(x)) = (-x)^2 = x^2.